OpenStudy (anonymous):

how do you do 4squareroot of 16x^8?

OpenStudy (anonymous):

\[2x^2\]

OpenStudy (anonymous):

i have the answer how did you solve it?

OpenStudy (anonymous):

oh it is \[\sqrt[4]{16x^8}\]

OpenStudy (anonymous):

first off \[\sqrt[4]{16}=2\] because \[2^4=16\]

OpenStudy (anonymous):

and also \[\sqrt[4]{x^8}=x^2\] because \[(x^2)^4=x^8\]

OpenStudy (anonymous):

which is another way of saying \[\frac{8}{4}=2\]

OpenStudy (anonymous):

thankyou!

OpenStudy (anonymous):

yw